if the ratio of the length to the width of a rectangle is 4:3 and the perimeter is 84 centimeters, find its area. Answer by dabanfield(803) 
 if the ratio of the length to the width of a rectangle is 4:3 and the perimeter is 84 centimeters, find its area.
Let 3x be the width and 4x the length. The ratio is then is 4x/3x = 4/3 or 4:3.
The perimeter 84 is equal to 2*w + 2*l = 2*3x + 2*4x or:
84 = 6x + 8x
14x = 84
x = 6
Now you can compute the width, 3x, and the length, 4x.